please help me solve this equation 0.1x-30=16-0.06x,80
Answer by stanbon(75887)

0.10x-30=16-0.06x
Add 0.06x to both sides to get; add 30 both sides to get:
0.16x = 46
x = 46/0.16
x = 287.50
=============
Cheers,
Stan H.